OpenStudy (jarynhoff):

How much interest will $2200 earn at 3% interest for 5 years, compounded annually? Give your answer in dollars and cents.

OpenStudy (calculusxy):

\[\large Y = a(b)^t\] a = principal amount b = rate t = time

OpenStudy (jarynhoff):

$333?

OpenStudy (calculusxy):

But remember for the rate, since this is an exponential growth question, you need to add a 1.

OpenStudy (calculusxy):

Let's substitute the amounts into its corresponding places. What is the principal (starting amount)?

OpenStudy (jarynhoff):

2200

OpenStudy (calculusxy):

Correct. Interest rate (as a decimal) ?

OpenStudy (jarynhoff):

.03

OpenStudy (calculusxy):

Yes and remember that it is important to add the 1 because its an exponential growth problem. So it would be 1 + 0.03 = 1.03 as \(b\).

OpenStudy (calculusxy):

Time?